# Accelerating Materials Design via LLM-Guided Evolutionary Search

### v1

@@ Abstract
Materials discovery requires navigating vast chemical and structural spaces while satisfying multiple, often conflicting, objectives. We present LLM-guided Evolution for MAterials design (LLEMA), a unified framework that couples the scientific knowledge embedded in large language models with chemistry-informed evolutionary rules and memory-based refinement. At each iteration, an LLM proposes crystallographically specified candidates under explicit property constraints; a surrogate-augmented oracle estimates physicochemical properties; and a multi-objective scorer updates success/failure memories to guide subsequent generations. Evaluated on 14 realistic tasks spanning electronics, energy, coatings, optics, and aerospace, LLEMA discovers candidates that are chemically plausible, thermodynamically stable, and property-aligned, achieving higher hit-rates and stronger Pareto fronts than generative and LLM-only baselines. Ablation studies confirm the importance of rule-guided generation, memory-based refinement, and surrogate prediction. By enforcing synthesizability and multi-objective trade-offs, LLEMA delivers a principled pathway to accelerate practical materials discovery.
Code: https://github.com/scientific-discovery/LLEMA
@@ 1 Introduction
Materials discovery requires identifying or designing materials with properties tailored to a specific task. However, the immense combinatorial space of chemical and structural compositions makes the traditional discovery process resource-intensive and slow (Hautier et al., 2012; Davies et al., 2016). While machine learning has accelerated the search for new materials, its reliance on large labeled datasets limits performance in data-scarce regimes. Trained on vast text corpora including scientific literature, Large language models (LLMs) offer a way to inject prior knowledge, making them tools for scientific discovery even in data-scarce settings (White, 2023). Recently, LLMs have been leveraged to bridge natural language and materials science, using textual knowledge to generate and refine hypotheses (Jia et al., 2024; Sprueill et al., 2024; Ghafarollahi & Buehler, 2025; Kumbhar et al., 2025). However, most existing methods rely on prompt engineering or unguided material generation, often producing candidates that are theoretically plausible yet unstable or impractical to synthesize. Moreover, they typically formulate material discovery as a single-objective task, optimizing for one property (e.g., bandgap, stability, or conductivity) in isolation, whereas real-world materials discovery is inherently multi-objective (see Figure 1), requiring trade-offs among competing targets such as electrical conductivity and thermal resistance in thermoelectric materials (Hao et al., 2019).
Figure 1: Overview of our multi-objective material discovery benchmark. The benchmark spans six diverse disciplines encompassing fourteen tasks with thermodynamic, electrical, physical, and chemical properties of materials.
@@
To address these challenges, we propose LLM-guided Evolution for MAterial discovery (LLEMA), a novel framework that uses LLM’s scientific knowledge with evolutionary search, and chemistry-informed design principles to generate and refine candidates under multiple task-specific property constraints. LLEMA introduces a range of chemistry-informed design principles as operators for candidate generation and refinement. These principles encode core knowledge across the entire materials discovery cycle, spanning compositional substitution, crystal structure manipulation, phase stability, and property-specific conditions. Unlike prior baselines, this chemically grounded generation integrates thermodynamic stability, enabling the systematic discovery of compounds that are both novel and experimentally realizable. At each iteration, the LLM fuses pretrained chemical knowledge with domain rules to balance exploration of chemical space and exploitation of promising leads. Candidates are expressed as crystallographic information files (CIFs) for downstream property prediction (Figure 2B). Surrogate-assisted oracle models then estimate task-relevant physicochemical properties (Figure 2C), and candidates are scored against both performance objectives and design constraints. Successful and failed trajectories are fed back to the LLM (Figure 2D), enabling iterative refinement of subsequent generations. Thus, LLEMA provides a principled framework for multi-objective discovery by explicitly enforcing stability and synthesizability to generate compounds that are not only novel but also physically realizable.
@@
To evaluate LLEMA, we introduce a benchmark suite of 14 diverse, industrially critical discovery tasks spanning wide-bandgap semiconductors, hard coatings, dielectrics, photovoltaics, and more (Figure 1). Unlike prior work that optimizes a single property in isolation, our tasks capture industrially relevant challenges that are inherently multi-objective. For example, discovering wide-bandgap semiconductors, a task critical for power and optoelectronic industries, requires simultaneously optimizing band gap and formation energy, rather than targeting either property alone. Such formulations reflect the reality of materials discovery, where progress hinges on balancing multiple constraints at once. The suite covers problems that are experimentally demanding and central to applications across electronics, energy, coatings, optics, and aerospace (Appendix B). By combining chemistry-informed design with iterative LLM-guided evolution, LLEMA goes beyond proposing candidates that are good at a single metric, instead generating plausible, synthesizable materials that satisfy complex, real-world objectives.
We evaluate LLEMA using GPT-4o-mini (OpenAI, 2023) and Mistral-Small-3.2-24B-Instruct-2506 (Jiang et al., 2024) as LLM backbones. Our results demonstrate that LLEMA consistently discovers chemically valid and structurally accurate candidates, with faster convergence across test settings. Our analysis further highlights the crucial role of rule-guided generation, memory-based feedback, and surrogate-assisted property prediction in LLEMA ’s performance. The major contributions of this work are as follows:
  • A unified, synthesizability-aware framework. We propose LLEMA, which integrates the scientific knowledge of LLMs with chemistry-informed evolutionary rules to generate candidates that are both novel and chemically synthesizable.
@@
  • Memory-based evolutionary refinement. We design a feedback mechanism that leverages success and failure pools, together with multi-island sampling, to iteratively steer LLM exploration toward high-performing regions while avoiding memorization.
  • Multi-objective discovery under realistic constraints. We formulate materials design as a constrained, multi-objective optimization problem and evaluate solutions via hit rate, thermodynamic stability, and Pareto front analysis.
  • Comprehensive evaluation across discovery tasks. We curate a benchmark suite of 14 industrially relevant discovery tasks, and demonstrate that LLEMA consistently discovers thermodynamically stable, chemically meaningful, and property-aligned candidates, outperforming prior baselines.
@@ 2.1 Problem Formulation
We formulate the material discovery task $\mathcal{T}$ as an optimization problem over the chemical space, where the goal is to identify the optimal material:
@@
where $m$ denotes a material from the valid candidate set $\mathcal{M}$ representing the chemical space, and the function $f:\mathcal{M}\rightarrow\mathbb{R}$ is a black-box objective that assigns each material a scalar value to the property of interest. However, in practice, material discovery rarely involves optimizing a single property. Instead, materials must satisfy multiple property constraints $\mathcal{C}=\{c_{1},c_{2},\dots,c_{k}\}$ while jointly optimizing competing objectives $f_{1},\dots,f_{n}$. Each constraint $c_{i}$ corresponds to a property function $f_{i}:\mathcal{M}\rightarrow\mathbb{R}$ and is satisfied when
$c_{i}:f_{i}(m)\in[l_{i},u_{i}]\quad\text{or}\quad c_{i}:f_{i}(m)\geq\tau_{i},$ (2)
where $l_{i},u_{i},\tau_{i}\in\mathbb{R}$ denote constraint-specific thresholds. The overall task thus reduces to identifying candidates $m\in\mathcal{M}$ that satisfy all constraints while achieving favorable trade-offs across objectives. A naive strategy for multi-objective optimization is to aggregate objectives via a weighted sum:
@@
where $w_{i}$ denotes the weight of the $i$-th objective. This formulation captures the multi-objective nature of material discovery, where competing property goals must be jointly optimized under domain-specific constraints. The task, therefore, reduces to efficiently identifying candidate materials that satisfy all property requirements while maximizing overall performance across objectives.
@@ 2.2 Hypothesis Generation
As illustrated in Figure 2, LLEMA begins with material candidate generation, followed by physicochemical property prediction, candidate evaluation, and evolutionary refinement. This stage combines the generative capabilities of LLMs with domain-guided constraints to synthesize chemically plausible hypotheses aligned with quantitative property targets. At each iteration $n$, the LLM $\pi_{\theta}$ samples a batch of $b$ candidate materials $\mathcal{M}^{b}$ from the prompt $\mathbf{p}_{n}$. Appendix D.1, Figure 7 details the construction of this prompt, composed of four components:
@@
  • Chemistry-Informed Design Principles: After the initial generation ($n=0$), the prompt incorporates chemistry-informed design rules $\mathcal{R}$ (e.g., same-group elemental substitutions, stoichiometry-preserving replacements). These rules act as operators that encode domain knowledge, guiding the search toward chemically meaningful regions of the space while maintaining enough flexibility to allow for novel discoveries. Appendix D.2 details the evolution rules.
@@ 2.3 Physicochemical Property Prediction
Following candidate generation, LLEMA estimates the physicochemical properties of the CIF representation of each material using a hierarchical prediction system. For a given candidate $\mathcal{M}_{j}$, the workflow first queries the reference model, which retrieves property values from curated experimental and computational databases like Materials Project (Jain et al., 2013) through exact or similarity-based matching. For out-of-distribution candidates, which lie outside the coverage of this reference model, LLEMA employs surrogate models such as CGCNN (Xie & Grossman, 2018) and ALIGNN (Choudhary & DeCost, 2021) to provide predictions. This yields a property vector ${f}(m)\in\mathbb{R}^{d}$, where each component corresponds to a physicochemical attribute of interest. The vector is subsequently evaluated by a multi-objective scoring function against the design constraints $\mathcal{C}$. Additional details on the implementation of these surrogate models are provided in Appendix D.2.
@@ 2.4 Fitness Assessment and Memory Management
Candidates are evaluated using a multi-objective scoring function that measures the degree of alignment between their predicted properties and the target design constraints $\mathcal{C}$. For each candidate material $\mathcal{M}_{j}$ generated at iteration $n$, the set of predicted properties is denoted by ${f_{i}(\mathcal{M}_{j})}_{i=1}^{k}$. The composite score is then computed as: $S({\mathcal{T,C;{\mathcal{M}}}_{j}})=\sum_{i=1}^{k}w_{i}\cdot\Phi_{i}(f_{i}(\mathcal{M}_{j}),c_{i})$; where $w_{i}$ represents the relative importance of the $i$-th property, $c_{i}$ denotes the corresponding target constraint, and $\Phi_{i}(\cdot,\cdot)$ is a normalized reward function that quantifies satisfaction of the constraint $c_{i}$ by the predicted value $f_{i}(\mathcal{M}_{j})$. Each candidate is then assigned to one of two memory pools: the success pool $\mathbb{M}^{+}$, containing candidates that satisfy all hard constraints (i.e., $\Phi_{i}\geq 0$ for all $i$), and the failure pool $\mathbb{M}^{-}$, containing candidates with negative composite scores. To progressively improve candidate quality and guide the search toward property-compliant regions, LLEMA employs a memory-based evolutionary refinement step inspired by island-model strategies (Romera-Paredes et al., 2024; Shojaee et al., 2024; Abhyankar et al., 2025). The candidate population is divided into $m$ independent islands containing success ($\mathbb{M}^{+}$) and failure memory ($\mathbb{M}^{-}$), each initialized with a copy of the initial exemplars. This structure supports parallel exploration, enabling different regions of chemical space to evolve independently and explore a range of candidates. At each iteration $n$, we first select one of the $m$ islands using Boltzmann sampling (De La Maza & Tidor, 1992), with a score-based probability of choosing a cluster $i$: $P_{i}=\frac{exp(s_{i}/\tau_{c})}{\sum_{i}exp(s_{i}/\tau_{c})}$, where $s_{i}$ denotes the mean score of the $i$-th cluster and $\tau_{c}$ is the temperature parameter. Within the chosen island, candidates are sampled from memory to construct the next prompt $\mathbf{p}_{n+1}$. Specifically, top-$k$ selection is applied to $\mathbb{M}^{+}$ and $\mathbb{M}^{-}$ to provide explicit demonstrations of high-scoring exemplars along with constraint violations. This mixture of successful and failed candidates, combined with domain-specific evolution rules $\mathcal{R}$, forms the in-context examples that guide the LLM in generating new candidates.
@@
As outlined in Algorithm 1, we initialize both the population of materials $\mathcal{P}_{0}$ and the candidate pool $\mathcal{M}$ that stores feasible solutions under the given design constraints $\mathcal{C}$. Subsequently, we construct a prompt by combining the task description with the associated physicochemical design constraints for generating candidate materials. At each iteration $n$, top-$k$ entries from prior iterations, along with pre-defined evolution rules $\mathcal{R}$ (e.g., stoichiometry, oxidation state, substitution rules) provided by domain scientists, are injected to ensure chemical and structural validity while exploring the candidate space. The LLM uses the prompt to generate crystallographic information file (CIF)–based material representations. These candidates are then evaluated using an oracle predictor to estimate their physicochemical properties. A scoring function assesses their performance relative to the target design objectives, partitioning candidates into success or failure pools based on whether they satisfy the constraints. The memory is updated accordingly, and a balanced sampling from both success and failure trajectories is used to provide feedback to the LLM, ensuring both exploitation of high-performing regions and exploration of underexplored design spaces. This iterative loop continues for $N$ rounds, after which the optimized candidate set $\mathbb{M}^{+}$ is returned.

### v2

# LLEMA: Evolutionary Search with LLMs for Multi-Objective Materials Discovery

@@ Abstract
Materials discovery requires navigating vast chemical and structural spaces while satisfying multiple, often conflicting, objectives. We present LLM-guided Evolution for MAterials discovery (LLEMA), a unified framework that couples the scientific knowledge embedded in large language models with chemistry-informed evolutionary rules and memory-based refinement. At each iteration, an LLM proposes crystallographically specified candidates under explicit property constraints; a surrogate-augmented oracle estimates physicochemical properties; and a multi-objective scorer updates success/failure memories to guide subsequent generations. Evaluated on 14 realistic tasks that span electronics, energy, coatings, optics, and aerospace, LLEMA discovers candidates that are chemically plausible, thermodynamically stable, and property-aligned, achieving higher hit rates and improved Pareto front quality relative to generative and LLM-only baselines. Ablation studies confirm the importance of rule-guided generation, memory-based refinement, and surrogate prediction. By enforcing synthesizability and multi-objective trade-offs, LLEMA provides a principled approach to accelerating practical materials discovery.
Code [Uncaptioned image]: https://github.com/scientific-discovery/LLEMA
Dataset [Uncaptioned image]: https://huggingface.co/datasets/nikhilsa/LLEMABench
@@ 1 Introduction
Materials discovery requires identifying or designing materials with properties tailored to a specific task. However, the immense combinatorial space of chemical and structural compositions makes the traditional discovery process resource-intensive and slow (Hautier et al., 2012; Davies et al., 2016). While machine learning has significantly accelerated materials screening, its effectiveness is often constrained by the availability of large, high-quality labeled datasets, limiting performance in data-scarce regimes (Chang et al., 2022; Xu et al., 2023). Trained on vast text corpora, including scientific literature, large language models (LLMs) offer a means to inject prior knowledge, making them promising tools for scientific discovery in data-scarce settings (White, 2023). Recently, LLMs have been leveraged to bridge natural language and material science, using textual knowledge to generate and refine hypotheses (Jia et al., 2024; Sprueill et al., 2024; Ghafarollahi and Buehler, 2025; Kumbhar et al., 2025). However, most existing methods rely on prompt engineering or unguided material generation, often producing candidates that are theoretically plausible yet unstable or impractical to synthesize. Moreover, they typically formulate materials discovery as a single-objective task, optimizing for one property (e.g., bandgap, stability, or conductivity) in isolation, whereas real-world materials discovery is inherently multi-objective (Gopakumar et al., 2018), requiring trade-offs among competing targets such as electrical conductivity and thermal resistance in thermoelectric materials (Hao et al., 2019).
@@
To address these challenges, we propose LLM-guided Evolution for MAterial discovery (LLEMA), a novel agentic framework that integrates LLM scientific priors with evolutionary search and chemistry-informed design principles to autonomously generate, evaluate, and refine candidate materials under multiple task-specific property constraints. LLEMA introduces a set of chemistry-informed design principles that act as operators to guide the LLM to generate and refine candidates iteratively. These principles encode core knowledge across the entire materials discovery cycle, spanning compositional substitution, crystal structure manipulation, phase stability, and property-specific conditions. Unlike prior baselines, this chemically grounded generation integrates thermodynamic stability, enabling the systematic discovery of compounds that are both novel and experimentally realizable. At each iteration, the LLM fuses pretrained chemical knowledge with domain rules to balance exploration of chemical space and exploitation of promising leads. The candidates are then expressed as crystallographic information files (CIFs) for downstream property prediction (Figure 1B). Surrogate-assisted oracle models then estimate task-relevant physicochemical properties (Figure 1C), and candidates are scored against both performance objectives and design constraints. Successful and failed trajectories are fed back to the LLM (Figure 1D), enabling evolutionary refinement of subsequent generations. Thus, LLEMA provides a principled framework for multi-objective discovery by explicitly enforcing stability and synthesizability to generate compounds that are not only novel but also physically realizable.
Table 1: Existing materials discovery frameworks.
Method Domain Multi-objective Rule-Guided Evolutionary
Knowledge Optimization Generation Refinement
CDVAE (Xie et al., 2022)
G-SchNet (Gebauer et al., 2019)
DiffCSP (Jiao et al., 2023)
MatterGen (Zeni et al., 2025)
LLMatDesign (Jia et al., 2024)
LLEMA (Ours)
Table 1 compares LLEMA with traditional and LLM-based materials discovery methods. Traditional generative methods require task-specific retraining that limits their generalizability across materials design problems. Furthermore, they lack the extensive prior knowledge embedded in LLMs and feedback-driven refinement mechanisms. Methods like LLMatDesign employ sequential, single-step modifications with self-reflection to optimize candidates for a single property. In contrast, LLEMA combines chemistry-informed evolutionary operators, surrogate-augmented oracle, multi-objective scoring mechanisms, and memory-based evolution to achieve feedback-driven materials design.
@@
To evaluate LLEMA, we introduce a new benchmark suite of 14 diverse, industrially relevant discovery tasks. The benchmark focuses on application-critical settings, covering problems relevant to electronics, energy, coatings, optics, and aerospace (Appendix C). Unlike existing benchmarks that focus on optimizing a single material property, each task in our suite is formulated as an explicit multi-objective discovery problem satisfying multiple properties. For example, discovering wide-bandgap semiconductors, a task critical for power and optoelectronic industries, requires simultaneously optimizing band gap and formation energy, rather than optimizing for a property in isolation. By capturing the complexity of real-world materials design, this benchmark provides a rigorous testbed for data-efficient, multi-objective discovery under physical and chemical constraints. Combining chemistry-informed design with iterative LLM-guided evolution, LLEMA goes beyond proposing candidates that are good at a single metric, instead generating plausible, synthesizable materials that satisfy complex, real-world objectives. We evaluated LLEMA using GPT-4o-mini (OpenAI, 2023) and Mistral-Small-3.2-24B-Instruct-2506 (Mistral AI, 2025) as LLM backbones. Our results demonstrate that LLEMA consistently discovers chemically valid and structurally accurate candidates, with faster convergence across test settings. Our analysis further highlights the crucial role of rule-guided generation, memory-based feedback, and surrogate-assisted property prediction in LLEMA ’s performance. The contributions of this work are summarized as:
  • A synthesizability-aware evolutionary framework. We introduce LLEMA, a framework that integrates LLMs’ scientific knowledge with chemistry-informed evolutionary operators, explicitly enforcing chemical validity and thermodynamic feasibility during the search process.
@@
  • Memory-based evolution. We design a mechanism that leverages success and failure pools, together with multi-island sampling, to iteratively steer LLMs toward high-performing regions while avoiding memorization.
  • Constrained multi-objective formulation for materials discovery. We cast materials design as a constrained multi-objective optimization problem, jointly optimizing competing property objectives.
  • Large-scale evaluation on realistic discovery tasks. We construct a benchmark suite of 14 industrially motivated materials discovery tasks and demonstrate that LLEMA consistently achieves higher validity, stability, and Pareto efficiency than state-of-the-art baselines.
@@ 2.1 Problem Formulation
We formulate the materials discovery task $\mathcal{T}$ as an optimization problem over the chemical space, where the goal is to identify the optimal material:
@@
where $m$ denotes a material from the valid candidate set $\mathcal{M}$ representing the chemical space, and the function $f:\mathcal{M}\rightarrow\mathbb{R}$ is a black-box objective that assigns each material a scalar value to the property of interest. However, in practice, materials discovery rarely involves optimizing a single property. Instead, materials must satisfy multiple property constraints $\mathcal{C}=\{c_{1},c_{2},\dots,c_{k}\}$ while jointly optimizing competing objectives $f_{1},\dots,f_{n}$. Each constraint $c_{i}$ corresponds to a property function $f_{i}:\mathcal{M}\rightarrow\mathbb{R}$ and takes one of the following canonical forms: an interval constraint, enforcing that the property lies within a feasible range; a lower-bound constraint, enforcing a minimum acceptable value; or an upper-bound constraint, enforcing a maximum allowable value. Formally,
$c_{i}:f_{i}(m)\in[l_{i},u_{i}]\quad\text{or}\quad c_{i}:f_{i}(m)\geq l_{i}\quad\text{or}\quad c_{i}:f_{i}(m)\leq u_{i},$ (2)
where $l_{i},u_{i}\in\mathbb{R}$ denote constraint-specific thresholds. The overall task thus reduces to identifying candidates $m\in\mathcal{M}$ that satisfy all constraints while achieving favorable trade-offs across objectives. A naive strategy for multi-objective optimization is to aggregate objectives via a weighted sum:
@@
where $w_{i}$ denotes the weight of the $i$-th objective. This formulation captures the multi-objective nature of materials discovery, where competing property goals must be jointly optimized under domain-specific constraints. The task, therefore, consists of efficiently identifying candidate materials that meet all property requirements while maximizing overall performance across objectives.
@@ 2.2 Hypothesis Generation
As illustrated in Figure 1, LLEMA begins with the generation of material candidates, followed by the prediction of physicochemical properties, the evaluation of candidates and the evolutionary refinement. This stage combines the generative capabilities of LLMs with domain-guided constraints to synthesize chemically plausible hypotheses aligned with quantitative property targets. At each iteration $n$, the LLM $\pi_{\theta}$ samples a batch of $b$ candidate materials $\mathcal{M}^{b}$ from the prompt $\mathbf{p}_{n}$. Appendix D.2, Figure 6 details the construction of this prompt, composed of four components:
@@
  • Chemistry-Informed Design Principles: After the initial generation ($n=0$), the prompt incorporates chemistry-informed design rules $\mathcal{R}$ (e.g., same-group elemental substitutions, stoichiometry-preserving replacements). These rules act as operators that encode domain knowledge, guiding the search toward chemically meaningful regions of the space while maintaining enough flexibility to allow for novel discoveries. Appendix B.4 details the evolution rules.
@@ 2.3 Physicochemical Property Prediction
Following candidate generation, LLEMA estimates the physicochemical properties of the CIF representation of each material using a hierarchical prediction system. For a given candidate $\mathcal{M}_{j}$, the workflow first queries the reference model, which retrieves property values from curated experimental and computational databases like Materials Project (Jain et al., 2013) through exact or similarity-based matching. For out-of-distribution candidates, which lie outside the coverage of this reference model, LLEMA employs surrogate models such as CGCNN (Xie and Grossman, 2018) and ALIGNN (Choudhary and DeCost, 2021) to provide predictions. This yields a property vector ${f}(m)\in\mathbb{R}^{d}$, where each component corresponds to a physicochemical attribute of interest. The vector is subsequently evaluated by a multi-objective scoring function against the design constraints $\mathcal{C}$. Additional details on the implementation of these surrogate models are provided in Appendix B.1.
@@ 2.4 Fitness Assessment and Memory Management
Candidates are evaluated using a multi-objective scoring function that measures the degree of alignment between their predicted properties and the target design constraints $\mathcal{C}$. For each candidate material $\mathcal{M}_{j}$ generated in iteration $n$, the set of predicted properties is denoted by ${f_{i}(\mathcal{M}_{j})}_{i=1}^{k}$. The composite score is then computed as
$S({\mathcal{T,C;{\mathcal{M}}}_{j}})=\sum_{i=1}^{k}w_{i}\cdot\Phi_{i}(f_{i}(\mathcal{M}_{j}),c_{i});$ (4)
where $w_{i}$ represents the relative importance of the $i$-th property, $c_{i}$ denotes the corresponding target constraint, and $\Phi_{i}(\cdot,\cdot)$ is a normalized reward function that quantifies the satisfaction of the constraint $c_{i}$ by the predicted value $f_{i}(\mathcal{M}_{j})$ (See Appendix B.2). Each candidate is then assigned to one of two memory pools: the success pool $\mathbb{M}^{+}$, containing candidates that satisfy all hard constraints (i.e., $S\geq 0$; $\Phi_{i}\geq 0$ for all $i$), and the failure pool $\mathbb{M}^{-}$, containing candidates that violate one or more constraints. To progressively improve candidate quality and guide the search toward property-compliant regions, LLEMA employs a memory-based evolutionary refinement step inspired by island-model strategies (Romera-Paredes et al., 2024; Shojaee et al., 2025a; Abhyankar et al., 2025). The candidate population is divided into $m$ independent islands containing success ($\mathbb{M}^{+}$) and failure memory ($\mathbb{M}^{-}$), each initialized with a copy of the initial exemplars. This structure supports parallel exploration, enabling different regions of the chemical space to evolve independently and explore a range of candidates. At each iteration $n$, we first select one of the $m$ ($m=5$) islands using Boltzmann sampling (De La Maza and Tidor, 1992), with a score-based probability of choosing a cluster $i$: $P_{i}=\frac{exp(s_{i}/\tau_{c})}{\sum_{j}exp(s_{j}/\tau_{c})}$, where $s_{i}$ denotes the mean score of the $i$-th cluster and $\tau_{c}$ is the temperature parameter. Within the chosen island, candidates are sampled from memory to construct the next prompt $\mathbf{p}_{n+1}$. Specifically, top-$k$ selection is applied to $\mathbb{M}^{+}$ and $\mathbb{M}^{-}$ to provide explicit demonstrations of high-scoring exemplars along with constraint violations. This mixture of successful and unsuccessful candidates, combined with domain-specific evolution rules $\mathcal{R}$, forms the in-context examples that guide the LLM in generating new candidates.
@@
As described in Algorithm 1, we initialize the population of materials $\mathcal{P}_{0}$ and the candidate pool $\mathcal{M}$ that stores feasible solutions under the given design constraints $\mathcal{C}$. Subsequently, we construct a prompt by combining the task description with the associated physicochemical design constraints for generating candidate materials. At each iteration $n$, top-$k$ entries from prior iterations (from one of the $m$ islands), along with pre-defined evolution rules $\mathcal{R}$ (e.g., stoichiometry, oxidation state, substitution rules) provided by domain scientists, are injected to ensure chemical and structural validity while exploring the candidate space. The LLM uses the prompt to generate crystallographic information file (CIF)–based material representations. These candidates are then evaluated using an oracle predictor to estimate their physicochemical properties. A scoring function assesses their performance relative to the target design objectives, partitioning candidates into success or failure pools based on whether they satisfy the constraints. Memory is updated accordingly and a balanced sampling from both success and failure trajectories is used to provide feedback to the LLM, ensuring both exploitation of high-performing regions and exploration of underexplored design spaces. This iterative loop continues for $N$ rounds, after which the optimized candidate set $\mathbb{M}^{+}$ is returned.